A Carnot heat engine works between 450oC and 25oC. The work output is used to drive...

  1. A Carnot heat engine works between 450oC and 25oC. The work output is used to drive a Carnot refrigerator that removes heat from -15oC space at a rate of 400kJ/min and rejects the heat to 25 oC environment.
  1. Find the rate of heat supplied to the heat engine
  2. Find the total rate of heat rejection to the environment

ii) In the steady state of a steam turbine operation, steam enters a turbine at a rate of 25000 kg/hr at 4MPa and 550oC. The steam expands and leaves the turbine as saturated vapor at 50kPa. The power produced by the turbine is 6MW.

  1. Given the surrounding is at 25oC, find the rate of entropy generation.

b. The turbine has an isentropic efficiency of 90%. Find the power output of the turbine.
Answer #1

The Combination of Heat Engine and Refrigrator is shown in the Diagram Above

COP=\frac{Q3}{Q4-Q3}COP=\frac{258}{298-258}=\frac{400}{Net Work}

Net Work = 62.01 KJ/Min

Efficiency of Engine =\frac{Net Work}{Qsupplied}Efficiency of Engine =\frac{723-523}{723}=\frac{62.01}{Qsupply}Qsupply= 224.186 KJ/min

Heat Rejected to Environment= Net work+Heat absorbed

Heat Rejected = 62.01 + 400

Heat Rejected = 462.015 KJ/min